Fincher ince has a total debit ratio of .64


2 value 1.00 points Problem 3-5 Calculating Leverage Ratios LO 2 Fincher, Inc., has a total debt ratio of 64 Wnat is its debt-equity ratio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g. 32.16.) Debl-equity ratio What is ts equity mutplier? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g, 32.1) Equity multiplier times round your answer to 2 decimal places 32 1 times Hints References eBook& Resources Hint
Answer #1

Answer 1:

Debt - equity ratio is ratio of total debt to shareholders' equity of a company.

Debt- equity ratio = Total debt / Shareholders' equity

Given:

Total debt = 0.64

Total debt ratio = Total liabilities / Total assets

Total assets = Total liabilities + Shareholders equity

When total debt ratio = 0.64

Equity proportion would be = 1 - 0.64 = 0.36

Debt- equity ratio = 0.64 / 0.36 = 1.78

1.78 times Debt - equity ratio

Answer 2:

Equity multiplier is the ratio of total assets divided by stockholder's equity. It is financial leverage ratio to evaluate company's use debt and equity in its capital structure to finance its assets.

When total debt ratio = 0.64

Equity proportion would be = 1 - 0.64 = 0.36

Equity multiplier = Total assets / Stockholders' equity = 1 /0.36 = 2.78

Equity multiplier 2.78 times